OpenStudy (umarsback):

Write the equation of a line in point-slope form that has a slope of 2/3 and passes through the point (2, -7).

OpenStudy (lucaz):

y-y'=m(x-x') y-(-7)=2(x-2)/3

OpenStudy (umarsback):

thank you!